Calculation to be done using RSuppose X has a standard normal distribution. The pnorm(), norm(), and qnorm() functions should be useful in the following. R
Answer #1

a)we have EaX +b) = aE(X) + b .Here E(X)=0

hence EC+1) = (E(X) + 1) (0 +1) =

b)we have VaX+b) = a V(X) .Here V(X)=1

Hence V*+1) = v(x + 1) = V(X) =

c)we have f(x) = be, -20 < x < 0

E(X) = L ***ax-v *e*ux

=2 | xe * V211 Jo(as the function is even)

=2 1 Gamma(+1) 2112since Xe-ar dir = m 1 Gamma(n+1) an

and Gamma(n)=(n-1)Gamma(n-1) and Gamma(1/2)=\sqrt{\Pi }

=1 gamma(7/2) 211 (1/2)7/ 2 1 5/2 * 3/2 * 1/2* VII 211 8vā - = 1

d)give the command >pnorm(0.83,mean=0,s.d=1)

it gives the probability that X<0.83

(1- above value) gives the required answer